What is the standard deviation of the returns on this stock?

State of the Economy        Probability       E(R)

Boom                                           0.33    24%

Normal 0.55    12%

Recession                                  0.12                      -60%
Answer #1

Expected return=Respective return*Respective probability

=(0.33*24)+(0.55*12)+(0.12*-60)=7.32%

probability Return probability*(Return-Expected Return)^2
0.33 24 0.33*(24-7.32)^2=91.813392
0.55 12 0.55*(12-7.32)^2=12.04632
0.12 -60 0.12*(-60-7.32)^2=543.837888
Total=647.6976%

Standard deviation=[Total probability*(Return-Expected Return)^2/Total probability]^(1/2)

=(647.6976)^(1/2)

=25.45%(Approx).